# Adaptive Local Combining with Decentralized Decoding for Distributed Massive MIMO

### v1

@@ IV-B Performance Evaluation
Figure 5: Distribution of Strong Pilot Decisions: How Each AP Adapts?
Fig. 5 shows the distribution of strong pilot decisions ($L_{\mathcal{S}_{m}}$) across all APs. The baseline PFZF and PWPFZF schemes exhibit a pronounced peak at $L_{\mathcal{S}_{m}}=6$-77, revealing their rigid threshold-based strategy that sacrifices excessive degrees of freedom for interference suppression regardless of local conditions. In striking contrast, the proposed G-PFZF and G-PWPFZF schemes show fundamentally different behavior: G-PFZF peaks at 2-4 strong pilots while G-PWPFZF shows an even more conservative distribution peaking at 1-3. This efficient allocation preserves spatial resources for signal enhancement. Most notably, the significant non-zero count at $L_{\mathcal{S}_{m}}=0$ demonstrates our algorithm’s ability to adaptively default to simple MR combining when it is the optimal strategy for a given AP’s local conditions. This demonstrates a seamless, adaptive switching between local zero-forcing and local MR processing that is impossible for the threshold-bound baseline.
@@
Fig. 6 depicts the sum spectral efficiency (SE) versus the number of UEs ($T$) for various combining schemes. The proposed distributed optimization framework delivers significant and scalable performance gains, with G-PFZF outperforming its baseline PFZF by 6.5% to 9.5% and G-PWPFZF surpassing PWPFZF by 5.5% to 10%, with the margin widening as the number of UEs increases. This superior performance highlights the critical limitation of the baseline PFZF and PWPFZF schemes, which rely on a hard, network-wide threshold using LSFCs to rigidly classify users as strong or weak at every AP. In contrast, the proposed framework introduces a distributed optimization algorithm, executed independently by each AP, to determine a better pilot-based grouping strategy. This allows an AP to adaptively select none, a subset, or all of the pilots to designate as strong, based on its local channel conditions and specific geometric configuration. This per-AP flexibility enables a more efficient and dynamic trade-off between utilizing degrees of freedom for interference suppression and desired signal enhancement. Consequently, the proposed G-PWPFZF scheme achieves superior performance over the G-PFZF scheme by suppressing interference from strong UEs for the weak UEs without a significant sacrifice in degrees of freedom.
@@
Fig. 7 illustrates the sum SE versus the number of orthogonal pilots $L_{p}$, demonstrating the superior performance of the proposed G-PFZF and G-PWPFZF schemes over their respective baseline counterparts. The proposed schemes achieve significant performance improvements of 6-14% for G-PFZF and 6-13% for G-PWPFZF compared to their respective baselines across different pilot lengths, with the performance gap widening as the number of pilots increases. This expanding margin highlights the advantage of the proposed adaptive user grouping strategy over the rigid network-wide threshold approach used in baseline schemes. The baseline schemes, constrained by a rigid network-wide threshold, must rigidly limit the number of strong pilots to avoid violating the fundamental condition for zero-forcing ($L_{\mathcal{S}_{m}}>A$), which leads to inefficient interference management. In contrast, our proposed distributed framework empowers each AP to independently and optimally select its set of strong pilots, inherently ensuring the selection is both locally optimal and always feasible ($L_{\mathcal{S}_{m}}<A$). This inherent adaptability allows our methods to leverage larger pilot sets more effectively, maximizing array gain and suppressing interference, which explains the observed widening of the performance gap. The G-PWPFZF scheme provides a further advantage by proactively protecting weak UEs from the strong UEs’ interference.
@@
Fig. 8 illustrates the sum spectral efficiency (SE) versus the number of antennas per AP ($A$) for various combining schemes, revealing a characteristic performance pattern that highlights the adaptability of our proposed framework. The G-PFZF and G-PWPFZF schemes demonstrate consistent superiority across all antenna configurations, with the performance gap exhibiting a distinctive trajectory that reflects adaptive resource utilization. At the minimal antenna configuration (2 antennas), where spatial degrees of freedom $A-L_{\mathcal{S}_{m}}$ are severely constrained, both proposed and baseline schemes face fundamental challenges in interference suppression. Nevertheless, our adaptive grouping strategy achieves measurable gains of 2.5% for G-PFZF and 6% for G-PWPFZF through an adaptive selection of combining schemes, dynamically switching between the G-PFZF/G-PWPFZF, and the MR at each AP based on local conditions. In the mid-range antenna configuration (4-8 antennas), where careful trade-off between interference suppression and signal enhancement becomes crucial, our proposed method demonstrates performance gain of 10-14% for G-PFZF and 10-18% for G-PWPFZF, showcasing their superior grouping capabilities. As antenna numbers increase further (10-16 antennas), providing abundant spatial resources that benefit all schemes, the performance gap narrows to 4.5-6.5% for G-PFZF and 3.5-5.5% for G-PWPFZF, though our approach maintains consistent superiority. These results confirm that our adaptive optimization framework achieves maximum relative advantage in precisely those scenarios where adaptive resource allocation is most valuable, particularly in resource-constrained environments, while still delivering meaningful gains in antenna-rich deployments.
@@
Fig. 9 illustrates the sum spectral efficiency (SE) versus the number of UEs ($T$), comparing local and LSFD-based combining schemes. For our G-PFZF and G-PWPFZF schemes, the performance gap between local weights and optimal LSFD remains minimal at just 3% across all UE densities. This marginal performance difference is dramatically outweighed by the substantial system benefits: our local weight approach eliminates the need for complex network-wide coordination, reducing computational cost at the CPU by avoiding large-scale matrix inversions and cutting fronthaul overhead by eliminating the need to share channel state information across the network. Despite this streamlined architecture, our locally-weighted G-PFZF and G-PWPFZF schemes still outperform the baseline PFZF and PWPFZF schemes with optimal LSFD by margins of 4.5-9% and 3.5-7.5% respectively. This demonstrates that the gains from our adaptive per-AP grouping strategy are so substantial that they outperform the benefits of optimal coordination applied to suboptimal threshold-based grouping. These results confirm that our fully distributed approach achieves an optimal balance between performance and practicality.
@@
Fig. 10 illustrates the sum spectral efficiency (SE) versus the number of pilots ($L_{p}$), comparing local and LSFD-based combining schemes. The performance gap between local and LSFD-based weights is less than 5.5% for sparse pilot configurations, where high pilot contamination makes centralized LSFD better equipped to handle interference. However, as the number of pilots increases to moderate and high values, the gap reduces to less than 1%, demonstrating near-identical performance. This trend occurs because with sufficient orthogonal pilots, both approaches have similar capability to mitigate pilot contamination. Notably, even in the worst-case scenario with limited pilots, the performance penalty remains modest, validating that the substantial fronthaul and complexity reductions of our decentralized decoding approach come with acceptable performance trade-offs across all pilot sequences. Furthermore, the proposed G-PFZF and G-PWPFZF schemes with local weights consistently outperform traditional PFZF and PWPFZF with LSFD across moderate to high pilot lengths, confirming the advantage of adaptive per-AP combining strategies.
@@
Fig. 11 presents the 90%-likely per-user spectral efficiency across different combining schemes, demonstrating remarkable improvements in user fairness through our distributed optimization framework. The proposed G-PFZF scheme achieves a substantial 45% improvement in 90%-likely SE compared to conventional PFZF, while G-PWPFZF shows a 34% gain over its baseline PWPFZF. These dramatic improvements confirm that our adaptive per-AP grouping strategy not only enhances overall system capacity but fundamentally transforms user experience across the network. The gains in 90%-likely SE particularly highlight our framework’s effectiveness in improving the performance of weak users who traditionally suffer from poor service quality. These results demonstrate that our distributed optimization framework successfully addresses both system-level efficiency and user-level fairness requirements in D-mMIMO deployments, achieving gains in quality-of-service uniformity.
@@
Fig. 12 presents the 90%-likely per-user spectral efficiency comparing local versus LSFD processing, revealing crucial insights about the performance-complexity trade-off in our distributed framework. While our G-PFZF and G-PWPFZF schemes with local weights experience a modest performance reduction of approximately 10-10.7% compared to their LSFD-based counterparts, the 10% performance difference between local and LSFD processing represents a reasonable trade-off for achieving full distributability, substantially reduced fronthaul overhead, and lower computational cost. Remarkably, our locally-weighted G-PFZF scheme outperforms the baseline PFZF with optimal LSFD by 32%, and G-PWPFZF with local weights surpasses the LSFD-enhanced PWPFZF baseline by 22.5%. These results demonstrate that the performance gains from our adaptive per-AP grouping strategy are so significant that they outweigh the benefits of optimal LSFD coordination applied to suboptimal threshold-based grouping. This analysis confirms that our distributed optimization framework provides a balance between performance and practical implementation constraints, delivering better SE per user while maintaining architectural advantages crucial for deployments.
V Conclusion
This paper has addressed the critical bottlenecks of sub-optimal performance, high fronthaul load, and computational cost in distributed massive MIMO networks by introducing a novel, decentralized decoding uplink architecture that fundamentally departs from conventional LSFD-based designs.
@@ V Conclusion
Through the proposed local ZF framework, we have demonstrated that enabling each AP to independently determine its combining strategy via local optimization, classifying pilots as strong or weak without fixed thresholds and dynamically switching among PFZF/PWPFZF, FZF, or MR, yields significant performance improvements. The resulting generalized schemes, G-PFZF and G-PWPFZF, consistently outperform their fixed-threshold counterparts across all evaluated scenarios, achieving substantial gains in both sum and per user spectral efficiency. Moreover, the introduction of pilot-dependent combining vectors and interference-suppressing local weights applied distributively at each AP eliminates the need for centralized LSFD, thereby drastically reducing fronthaul overhead and computational cost. Remarkably, this distributed approach incurs only a minimal performance penalty compared to idealized centralized coordination, while even outperforming conventional threshold-based schemes using optimal LSFD.
These findings collectively establish that adaptive processing at the network, not complex centralized decoding, is the key to scalability and efficiency in future D-mMIMO systems. The proposed framework offers a practical and high-performance pathway toward realizing scalable distributed MIMO networks without compromising on quality-of-service or imposing prohibitive infrastructure costs.
-A Proof of closed-form SINR expression for G-PFZF combining:
@@
To derive the closed-form expression, we need few expectation properties,
$\begin{split}\mathbb{E}\Big\{\Big|\sum_{m}a_{m}\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big|^{2}\Big\}=\sum_{m}|a_{m}|^{2}\mathbb{E}\Big\{\Big|\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big|^{2}\Big\}+\Big|\sum_{m}a_{m}\mathbb{E}\Big\{\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big\}\Big|^{2}-\sum_{m}\Big|a_{m}\mathbb{E}\Big\{\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big\}\Big|^{2},\end{split}$ (35)
Also, if $\mathbf{v}_{m}$ and $\mathbf{g}_{m}$ are indpendent random vectors of N length with zero mean and elements following i.i.d., then
$\mathbb{E}\Big\{\Big|\mathbf{v}_{m}^{H}\mathbf{g}_{m}\Big|^{2}\Big\}=\frac{\mathbb{E}\Big\{\Big|\mathbf{v}_{m}\Big|^{2}\Big\}\mathbb{E}\Big\{\Big|\mathbf{g}_{m}\Big|^{2}\Big\}}{N}.$ (36)
We derive the closed-form expression of desired signal term in (6) using the combining vectors in (7) and (8).
$|\text{DS}_{t}|^{2}=p^{u}_{t}\Big|\mathbb{E}\Big\{\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mt}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mt}\Big\}\Big|^{2},$ (37)
@@
where $\mathcal{Z}_{i_{t}}$ is the subset of APs for which pilot $i_{t}$ is strong and $\mathcal{Y}_{i_{t}}$ is the subset of APs for which pilot $i_{t}$ is weak. Using (10) and (12), the (37) can be written as:
$|\text{DS}_{t}|^{2}=p^{u}_{t}\Bigg|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}\sqrt{(A-L_{\mathcal{S}_{m}})\gamma_{mk}}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}\sqrt{A\gamma_{mk}}\Bigg|^{2}=p^{u}_{t}\Bigg|\sum_{m=1}^{M}a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Bigg|^{2},$ (38)
where $\delta_{mi_{t}}$ is defined in (15). The first term of interference in (6) can be expanded as:
$\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}=p^{u}_{t}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mt}+\sum_{m\in\mathcal{Y}_{i_{t}}}a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mt}\Big|^{2}\Bigg\}-|\text{DS}_{t}|^{2},$ (39)
We focus on first term of $\mathbb{E}\Big\{|\text{BU}_{t}|^{2}\Big\}$, which can be written as:

### v2

@@ IV-B Performance Evaluation
Figure 3: Distribution of strong pilot decisions.
Fig. 3 shows the distribution of strong pilot decisions ($L_{\mathcal{S}_{m}}$) across all APs. The baseline PFZF and PWPFZF schemes exhibit a pronounced peak at $L_{\mathcal{S}_{m}}{=}6$-77, revealing their rigid threshold-based strategy that sacrifices excessive degrees of freedom for interference suppression regardless of local conditions. In striking contrast, the proposed G-PFZF and G-PWPFZF schemes show fundamentally different behavior: G-PFZF peaks at 2-4 strong pilots while G-PWPFZF shows an even more conservative distribution peaking at 1-3. This efficient allocation preserves spatial resources for signal enhancement. Most notably, the significant non-zero count at $L_{\mathcal{S}_{m}}{=}0$ demonstrates our algorithm’s ability to adaptively default to simple MR combining when it is the optimal strategy for a given AP’s local conditions. This demonstrates a seamless, adaptive switching between local zero-forcing and local MR processing that is impossible for the threshold-bound baselines.
@@
Fig. 4 depicts the sum SE versus the number of UEs ($T$) for various combining schemes. The proposed decentralized optimization framework delivers significant and scalable performance gains, with G-PFZF outperforming its baseline PFZF by 6.5% to 9.5% and G-PWPFZF surpassing PWPFZF by 5.5% to 10%, with the gaps widening as the number of UEs increases. Better performance of the proposed schemes highlights the critical limitation of the baseline PFZF and PWPFZF schemes, which rely on a hard, network-wide threshold using LSFCs to rigidly classify users as strong or weak at every AP. In contrast, the proposed framework introduces a decentralized optimization algorithm, executed independently by each AP, to determine a better pilot-based grouping strategy. This allows an AP to adaptively select none, a subset, or all of the pilots as strong, based on its local channel conditions and specific geometric configuration. This per-AP flexibility enables a more efficient and dynamic trade-off between utilizing degrees of freedom for interference suppression and desired signal enhancement. Consequently, the proposed G-PWPFZF scheme achieves better performance over the G-PFZF scheme by suppressing interference of the weak UEs from the strong UEs at the APs.
@@
Fig. 5 illustrates the sum SE versus the number of orthogonal pilots $L_{p}$, demonstrating the improved performance of the proposed G-PFZF and G-PWPFZF schemes over their respective baseline counterparts. The proposed schemes achieve significant performance improvements of 6-14% for G-PFZF and 6-13% for G-PWPFZF compared to their respective baselines across different pilot lengths, with the performance gap widening as the number of pilots increases. This expanding gap highlights the advantage of the proposed adaptive user grouping strategy over the rigid network-wide threshold approach used in baseline schemes. The baseline schemes, must rigidly limit the number of strong pilots to avoid violating the fundamental condition for zero-forcing ($L_{\mathcal{S}_{m}}{>}A$), which leads to inefficient interference management. In contrast, our proposed decentralized framework empowers each AP to independently and optimally select its set of strong pilots, inherently ensuring that the selection is both locally optimal and always feasible ($L_{\mathcal{S}_{m}}{<}A$). This inherent adaptability allows our methods to leverage larger pilot sets more effectively, maximizing array gain and suppressing interference, which explains the observed widening of the performance gap with the number of pilots.
@@
Fig. 6 illustrates the sum SE versus the number of antennas per AP ($A$) for various combining schemes, revealing a characteristic performance pattern that highlights the adaptability of our proposed framework. The G-PFZF and G-PWPFZF schemes demonstrate consistent superiority across all antenna configurations, with the performance gap exhibiting a distinctive trajectory that reflects adaptive resource utilization. At the minimal antenna configuration (2 antennas), where spatial degrees of freedom $A-L_{\mathcal{S}_{m}}$ are severely constrained, both proposed and baseline schemes face fundamental challenges in interference suppression. Nevertheless, our adaptive grouping strategy achieves measurable gains of 2.5% for G-PFZF and 6% for G-PWPFZF through an adaptive selection of combining schemes, dynamically switching between the G-PFZF/G-PWPFZF, and the MR at each AP based on local conditions. In the mid-range antenna configuration (4-8 antennas), where careful trade-off between interference suppression and signal enhancement becomes crucial, our proposed method demonstrates performance gain of 10-14% for G-PFZF and 10-18% for G-PWPFZF, showcasing their superior grouping capabilities. As antenna numbers increase further (10-16 antennas), providing abundant spatial resources that benefit all schemes, the performance gap narrows to 4.5-6.5% for G-PFZF and 3.5-5.5% for G-PWPFZF, though our approach maintains consistent superiority.
@@
Fig. 7 illustrates the sum SE versus the number of UEs ($T$), comparing local and LSFD-based combining schemes. For our G-PFZF and G-PWPFZF schemes, the performance gap between local weights and optimal LSFD remains minimal at just 3% across all UE densities. This marginal performance difference is dramatically outweighed by the substantial system benefits: our local weight approach eliminates the need for complex network-wide coordination, reducing computational cost at the CPU by avoiding large-scale matrix inversions and reducing fronthaul overhead by eliminating the need to share channel state information across the network. Despite this streamlined architecture, our locally-weighted G-PFZF and G-PWPFZF schemes still outperform the baseline PFZF and PWPFZF schemes with LSFD by margins of 4.5-9% and 3.5-7.5% respectively. This demonstrates that the gains from our adaptive per-AP grouping strategy are so substantial that they outperform the benefits of optimal coordination applied to suboptimal threshold-based grouping. These results confirm that our fully decentralized approach achieves a better balance between performance and practicality.
@@
Fig. 8 illustrates the sum SE versus the number of pilots ($L_{p}$), comparing local and LSFD-based combining schemes. The performance gap between local and LSFD-based weights is less than 5.5% for sparse pilot configurations, where high pilot contamination makes centralized LSFD better equipped to handle interference. However, as the number of pilots increases to moderate and high values, the gap reduces to less than 1%, demonstrating near-identical performance. This trend occurs because with sufficient orthogonal pilots, both approaches have similar capability to mitigate pilot contamination. Notably, even in the worst-case scenario with limited pilots, the performance penalty remains modest, validating that the substantial fronthaul and complexity reductions of our decentralized decoding approach come with acceptable performance trade-offs across all pilot sequences. Furthermore, the proposed G-PFZF and G-PWPFZF schemes with local weights consistently outperform PFZF and PWPFZF with LSFD across moderate to high pilot lengths, confirming the advantage of adaptive per-AP combining strategies.
@@
Fig. 9 presents the 90%-likely per-user SE across different combining schemes, demonstrating significant improvements in user fairness through our decentralized optimization framework. The proposed G-PFZF scheme achieves 45% improvement in 90%-likely SE compared to conventional PFZF, while G-PWPFZF shows a 34% gain over its baseline PWPFZF. These gains confirm that our adaptive per-AP grouping strategy not only enhances the overall system capacity but fundamentally transforms user experience across the network. The gains in 90%-likely SE particularly highlight our framework’s effectiveness in improving the performance of weak users.
Figure 10: The uplink SE per user comparison for various decoding schemes.
@@
Fig. 10 presents the 90%-likely per-user SE comparing local versus LSFD processing, revealing crucial insights about the performance-complexity trade-off in our decentralized framework. While our G-PFZF and G-PWPFZF schemes with local weights experience a modest performance reduction of approximately 10-10.7% compared to their LSFD-based counterparts, the 10% performance difference between local and LSFD processing represents a reasonable trade-off for achieving fully decentralized architecture, substantially reduced fronthaul overhead, and lower computational cost. Noticeably, our locally-weighted G-PFZF scheme outperforms the baseline PFZF with optimal LSFD by 32%, and G-PWPFZF with local weights surpasses the LSFD-enhanced PWPFZF baseline by 22.5%. These results demonstrate that the performance gains from our adaptive per-AP grouping strategy are so significant that they outweigh the benefits of optimal LSFD coordination applied to suboptimal threshold-based grouping.
V Conclusion and Future Work
@@
This paper has addressed the critical bottlenecks of sub-optimal performance, high fronthaul load, and computational cost in D-mMIMO networks by introducing a novel, decentralized decoding uplink architecture that fundamentally departs from conventional LSFD-based designs. Through the proposed local ZF framework, we have demonstrated that enabling each AP to independently determine its combining strategy via local optimization, classifying pilots as strong or weak without fixed thresholds and dynamically switching among PFZF/PWPFZF, FZF, or MR, yields significant performance improvements. The resulting generalized schemes, G-PFZF and G-PWPFZF, consistently outperform their fixed-threshold counterparts across all evaluated scenarios, achieving substantial gains in both sum and per user spectral efficiency. Moreover, the introduction of pilot-dependent combining vectors and interference-suppressing local weights applied distributively at each AP eliminates the need for centralized LSFD, thereby drastically reducing fronthaul overhead and computational cost. Remarkably, this decentralized approach incurs only a small performance penalty compared to idealized centralized coordination, while even outperforming conventional threshold-based schemes using optimal LSFD. These findings collectively establish that proposed decentralized decoding in the network, not centralized decoding, is the key to scalability and efficiency in future D-mMIMO systems without compromising on quality-of-service.
In future, we plan to investigate the robustness of the proposed decentralized architecture and generalized combining schemes under more practical and challenging conditions, such as hardware impairments.
-A Derivation of closed-form SINR expression for G-PFZF combining
@@
In order to derive the closed-form expression, we need few expectations computation or expansions
$\begin{split}\mathbb{E}\Big\{\Big|\!\!\sum_{m}\!\!\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\!\Big|^{2}\!\Big\}\!\!=&\!\sum_{m}\!\mathbb{E}\Big\{\!\Big|\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\!\Big|^{2}\!\Big\}+\!\Big|\!\sum_{m}\!\mathbb{E}\Big\{\mathbf{v}_{mi}^{H}\mathbf{g}_{m}\Big\}\Big|^{2}-\sum_{m}\Big|\mathbb{E}\Big\{\mathbf{v}_{mi}^{H}\mathbf{g}_{mi}\Big\}\Big|^{2},\end{split}$ (A1)
$\mathbb{E}\Big\{\Big|\mathbf{v}_{mi_{t}}^{H}\tilde{\mathbf{g}}_{mk}\Big|^{2}\Big\}=$ $\beta_{mk}-\gamma_{mk}$ (A2)
$\mathbb{E}\Big\{\Big|\mathbf{v}_{mi_{t}}^{H}{\mathbf{n}}_{m}\Big|^{2}\Big\}=$ $1.$ (A3)
The desired signal term in (6) can be written as
$|\text{DS}_{t}|^{2}=\Big|\mathbb{E}\Big\{\!\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\hat{\mathbf{{g}}}_{mt}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\hat{\mathbf{{g}}}_{mt}\Big\}\Big|^{2},$
@@
where $\mathcal{Z}_{i_{t}}$ and $\mathcal{Y}_{i_{t}}$ denote the subsets of APs where pilot $i_{t}$ is classified as strong and weak, respectively. Using (10) and (12), the $|\text{DS}_{t}|^{2}$ can be written as:
$\begin{split}|\text{DS}_{t}|^{2}&=\!\!\Big|\!\!\!\!\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{(A{-}L_{\mathcal{S}_{m}})\gamma_{mk}}+\!\!\!\!\!\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}\sqrt{A\gamma_{mk}}\Big|^{2}\\ &=\Big|\sum_{m=1}^{M}\!a_{mt}\sqrt{(A-\delta_{mi_{t}}L_{\mathcal{S}_{m}})\gamma_{mk}}\Big|^{2},\end{split}$ (A4)
The interference term $\mathbb{E}\big\{|\text{UI}_{tk}|^{2}\big\}$ in (6) can be expanded as:
$\begin{split}\mathbb{E}\Bigg\{\Big|\sum_{m\in\mathcal{Z}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\mathbf{{g}}_{mk}+\sum_{m\in\mathcal{Y}_{i_{t}}}\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\mathbf{{g}}_{mk}\Big|^{2}\Bigg\}&=\!\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\!\Bigg\}\!+\mathbb{E}\Bigg\{\!\Big|\!\!\!\!\!\!\!\sum_{\ \ \ m\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!\!\!a_{mt}(\mathbf{v}_{mi_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{mk}\Big|^{2}\!\!\Bigg\}\\ &+2\mathbb{E}\Bigg\{\!\!\!\!\sum_{\ \ \ m\in\mathcal{Z}_{i_{t}}}\!\!\!\!\!\!\!\!\sum_{\ \ \ n\in\mathcal{Y}_{i_{t}}}\!\!\!\!\!\!a_{mt}a_{nt}(\mathbf{v}_{mi_{t}}^{\text{LZF}})^{H}\!\mathbf{{g}}_{mk}(\mathbf{v}_{ni_{t}}^{\text{MR}})^{H}\!\mathbf{{g}}_{nk}\!\Bigg\}.\end{split}$ (A5)
Using (A1), the first term ($A_{1}$) of (A5) can be rewritten as: